Grade 9Magical Mathematics[Interesting Approach]

A,B and C start moving from a point at the same time in circle....A and B clocwise.....and C anticlocwise.....if A meets C in 88secs....B meets C in 110secs..In what time will A meet B PLS SHOW THE STEPS!!!!

Shivam Bhagat
15 Years agoGrade 9

1 Answer

andrew hunt potter
15 Years ago

w=angular valocity

w(a/c)=wa-(-wc )=wa+wc       ...........................(as c moves in opposite direction)       

2π/(wa+wc )=88

wa+wc =2π/88.............1

similarly,wb+wc =2π/110...............2

subtracting eqn 1 & 2

wa-wb=2π/(1/88-1/110)=π/220

wa/b= π/220

in b frame b stationery.... a move & meet by movin a 360 degrees ;

time= 2π/wa/b =440 seconds